(3 points) A careless university student leaves her iClicker device behind with probability 1/4 each time she attends a class. She sets out with her iClicker device to attend 5 different classes (each class is in a different lecture theatre). Part 1) If she arrives home without her iClicker device (after attending 5 classes), what is the probability (to 3 SIGNIFICANT figures) that she left it in the 5th class? Probability0.079 Part 2) If she arrives home without her iClicker device and she is sure she has the iClicker device after leaving the first class, what is the probability (to 3 SIGNIFICANT figures) that she left it in the 5th class? Probability0.105
Answer #1

1)P(lost in 5th class |i clicker lost) =P(not in first 4 but in 5th)/P(lost)

=((1-0.25)4*0.25)/(1-0.755) =0.1037

2)

P(lost in 5th class |i clicker lost but not in 1st) =((1-0.25)3*0.25)/(1-0.754)=0.1543